# Price predictability at ultra-high frequency: Entropy-based randomness test

### v2

@@ Appendix B Statistics and partitions of predictable days
parameter
mean for predictable days
mean for not predictable days
p-value
Sample size (number of days)
14
66
number of non-zero returns
2685
2454
0.177
fraction of 0-returns
0.100
0.094
0.143
$k$
6
5.758
7.594 ×$\times 10^{-5}$**
$(\hat{p}(0\dots 0)+\hat{p}(1\dots 1))2^{k}$
3.033
2.462
0.013*
$|\hat{p}(1)-\hat{p}(0)|$
0.068
0.039
0.003**
magnitude of daily log-price increment
0.015
0.014
0.952
mean price returns
-3.124×$\times 10^{-7}$
-6.043×$\times 10^{-7}$
0.874
magnitude of autocorrelation of non-zero returns
0.0299
0.0297
0.966
magnitude of autocorrelation of absolute values
0.079
0.093
0.208
$\nu$ of t-distribution
2.656
3.426
0.435
scale of t-distribution
1.843×$\times 10^{-4}$
2.156×$\times 10^{-4}$
0.04*
magnitude of shift of t-distribution
1.609×$\times 10^{-5}$
1.204×$\times 10^{-5}$
0.11
daily volume
12,788,633
12,048,587
0.438
fraction of jumps
2.093×$\times 10^{-4}$
1.685×$\times 10^{-4}$
0.682
@@
parameter
mean for predictable days
mean for not predictable days
p-value
Sample size (number of days)
38
42
number of non-zero returns
2626
2195
0.018*
fraction of 0-returns
0.695
0.660
0.0007**
$k$
5.816
5.595
0.04*
$(\hat{p}(0\dots 0)+\hat{p}(1\dots 1))2^{k}$
2.5
2.25
0.118
$|\hat{p}(1)-\hat{p}(0)|$
0.0205
0.0203
0.956
magnitude of daily log-price increment
0.037
0.027
0.1
mean price returns
1.735×$\times 10^{-6}$
-1.491×$\times 10^{-6}$
0.016*
magnitude of autocorrelation of non-zero returns
0.132
0.081
0.001**
magnitude of autocorrelation of absolute values
0.286
0.234
0.001**
$\nu$ of t-distribution
1.816
1.356
0.002**
scale of t-distribution
1.281×$\times 10^{-5}$
2.868×$\times 10^{-5}$
0.226
magnitude of shift of t-distribution
1.672×$\times 10^{-7}$
4.625×$\times 10^{-7}$
0.194
daily volume
7,087,714
4,776,277
0.0001**
fraction of jumps
0.018
0.016
0.049*
@@
parameter
mean for predictable days
mean for not predictable days
p-value
Sample size (number of days)
29
51
number of non-zero returns
12764
10778
0.04*
fraction of 0-returns
0.170
0.168
0.669
$k$
6.862
6.745
0.217
$(\hat{p}(0\dots 0)+\hat{p}(1\dots 1))2^{k}$
2.881
2.539
8.075×$\times 10^{-5}$**
$|\hat{p}(1)-\hat{p}(0)|$
0.015
0.013
0.417
magnitude of daily log-price increment
0.011
0.008
0.214
mean price returns
-1.464×$\times 10^{-7}$
1.080×$\times 10^{-7}$
0.196
magnitude of autocorrelation of non-zero returns
0.026
0.018
0.009**
magnitude of autocorrelation of absolute values
0.086
0.089
0.76
$\nu$ of t-distribution
1.988
2.245
0.031*
scale of t-distribution
5.531×$\times 10^{-5}$
5.467×$\times 10^{-5}$
0.792
magnitude of shift of t-distribution
9.444×$\times 10^{-7}$
8.551×$\times 10^{-7}$
0.544
daily volume
10,155,452
8,554,678
0.037*
fraction of jumps
4.187×$\times 10^{-4}$
5.052×$\times 10^{-4}$
0.147

### v3

@@ Appendix B Statistics and partitions of predictable days
parameter mean for predictable days mean for not predictable days p-value
Sample size (number of days) 14 66
number of non-zero returns 2685 2454 0.177
fraction of 0-returns 0.100 0.094 0.143
$k$ 6 5.758 7.594 ×$\times 10^{-5}$**
$(\hat{p}(0\dots 0)+\hat{p}(1\dots 1))2^{k}$ 3.033 2.462 0.013*
$|\hat{p}(1)-\hat{p}(0)|$ 0.068 0.039 0.003**
magnitude of daily log-price increment 0.015 0.014 0.952
mean price returns -3.124×$\times 10^{-7}$6.043×$\times 10^{-7}$ 0.874
magnitude of autocorrelation of non-zero returns 0.0299 0.0297 0.966
magnitude of autocorrelation of absolute values 0.079 0.093 0.208
$\nu$ of t-distribution 2.656 3.426 0.435
scale of t-distribution 1.843×$\times 10^{-4}$ 2.156×$\times 10^{-4}$ 0.04*
magnitude of shift of t-distribution 1.609×$\times 10^{-5}$ 1.204×$\times 10^{-5}$ 0.11
daily volume 12,788,633 12,048,587 0.438
fraction of jumps 2.093×$\times 10^{-4}$ 1.685×$\times 10^{-4}$ 0.682
@@
parameter mean for predictable days mean for not predictable days p-value
Sample size (number of days) 38 42
number of non-zero returns 2626 2195 0.018*
fraction of 0-returns 0.695 0.660 0.0007**
$k$ 5.816 5.595 0.04*
$(\hat{p}(0\dots 0)+\hat{p}(1\dots 1))2^{k}$ 2.5 2.25 0.118
$|\hat{p}(1)-\hat{p}(0)|$ 0.0205 0.0203 0.956
magnitude of daily log-price increment 0.037 0.027 0.1
mean price returns 1.735×$\times 10^{-6}$1.491×$\times 10^{-6}$ 0.016*
magnitude of autocorrelation of non-zero returns 0.132 0.081 0.001**
magnitude of autocorrelation of absolute values 0.286 0.234 0.001**
$\nu$ of t-distribution 1.816 1.356 0.002**
scale of t-distribution 1.281×$\times 10^{-5}$ 2.868×$\times 10^{-5}$ 0.226
magnitude of shift of t-distribution 1.672×$\times 10^{-7}$ 4.625×$\times 10^{-7}$ 0.194
daily volume 7,087,714 4,776,277 0.0001**
fraction of jumps 0.018 0.016 0.049*
@@
parameter mean for predictable days mean for not predictable days p-value
Sample size (number of days) 29 51
number of non-zero returns 12764 10778 0.04*
fraction of 0-returns 0.170 0.168 0.669
$k$ 6.862 6.745 0.217
$(\hat{p}(0\dots 0)+\hat{p}(1\dots 1))2^{k}$ 2.881 2.539 8.075×$\times 10^{-5}$**
$|\hat{p}(1)-\hat{p}(0)|$ 0.015 0.013 0.417
magnitude of daily log-price increment 0.011 0.008 0.214
mean price returns -1.464×$\times 10^{-7}$ 1.080×$\times 10^{-7}$ 0.196
magnitude of autocorrelation of non-zero returns 0.026 0.018 0.009**
magnitude of autocorrelation of absolute values 0.086 0.089 0.76
$\nu$ of t-distribution 1.988 2.245 0.031*
scale of t-distribution 5.531×$\times 10^{-5}$ 5.467×$\times 10^{-5}$ 0.792
magnitude of shift of t-distribution 9.444×$\times 10^{-7}$ 8.551×$\times 10^{-7}$ 0.544
daily volume 10,155,452 8,554,678 0.037*
fraction of jumps 4.187×$\times 10^{-4}$ 5.052×$\times 10^{-4}$ 0.147